OpenStudy (anonymous):

(Trig) Find circumference of circle, given that length of arc (s) = 25km, angle of arc (O) = 20 degrees

OpenStudy (anonymous):

I know that lenght of arc/circumfrence = Angle of arc/360 degrees

OpenStudy (anonymous):

25/C = 20 degrees/ 360 degrees

OpenStudy (anonymous):

So first I isolated C, by multiplying both sides by 25

OpenStudy (anonymous):

that leaves me at C=20x25/360... but i think this is where i went wrong because when i simply this it gives me C=1.38, but the answer is 450km

OpenStudy (anonymous):

I know the answer but not sure how to get to it myself.

OpenStudy (anonymous):

Am I missing something here?

OpenStudy (anonymous):

circumference will be 25 * 18 = 450km

OpenStudy (anonymous):

\[\frac{25}{c}=\frac{20}{360}\]\[c=\frac{360 \times 25}{20}\]

OpenStudy (anonymous):

\[\frac{25}{c} = \frac{20}{360}\]times both sides by c\[25 = \frac{20c}{360}\]times both sides by 360\[25 \times 360 = 20c\]divide both sides by 20\[c=\frac{360 \times 25}{20}\]

OpenStudy (anonymous):

Thankyou, nphuongsun93, I appreciate your help.